When jumping straight down, you can be seriously injured if you land stiff-legged. One way to avoid injury is to bend your knees upon landing to reduce the force of the impact. A 75-kg man just before contact with the ground has a speed of 5.5 m/s.

(a) In a stiff-legged landing he comes to a halt in 2.4 ms. Find the average net force that acts on him during this time. _________N

(b) When he bends his knees, he comes to a halt in 0.12 s. Find the average force now. _______N

(c) During the landing, the force of the ground on the man points upward, while the force due to gravity points downward. The average net force acting on the man includes both of these forces. Taking into account the directions of these forces, find the force of the ground on the man in parts (a) and (b). stiff legged landing _________N bent legged landing ________N

Step by steps are appreciated, thanks in advance for your help!
Answer #1

here,

mass , m = 75 kg

initial speed , u = 5.5 m/s

a)

time taken , t = 2.4 ms = 0.0024 s

the average net force that acts on him during this time , F1= impulse /time

F1 = - m * u /t = - 75 * 5.5 /0.0024

F1= - 1.71875 * 10^5 N

b)

time taken , t = 0.12 s

the average net force that acts on him during this time , F2 = impulse /time

F2 = - m * u /t = - 75 * 5.5 /0.12

F2= - 3.4375 * 10^3 N

c)

for stiff legged landing ,

gravity is very smaller than F1 so,

the force of the ground on the man is 1.71875 * 10^5 N

fir bent legged landing , N2 = F2 + m * g

N2 = 3437.5 + 75 * 9.8 N

N2= 4.1725 * 10^3 N

feel free to drop comment if you have any doubt.